OpenStudy (anonymous):

what is the ordered pair solution to 0.5x+0.2y=89 0.4-0.5y=-8

OpenStudy (anonymous):

Is that 0.4x-0.5y = -8 or is it actually what you wrote?

OpenStudy (anonymous):

the two equations are supposedly able to solve using technology. but i dont know how. i wrote them exactly as they were given

OpenStudy (anonymous):

If you rearrange the second equation, you get: -0.5y = -8 - 0.4 y = (-8-0.4)/-0.5 y = -8.4 / - 0.5 y = 16.8 So we solved for y by rearranging and got y = 16.8. Now that we know what y is, we can plug in this value in to the first equation to solve for x. This is what it looks like: 0.5x + 0.2(16.8) = 89 0.5x = 89 - 0.2(16.8) x = [89 - 0.2(16.8)] / 0.5 x = [89 - 3.36] / 0.5 x = 85.64 / 0.5 x = 171.28 Therefore, x = 171.28, and y = 16.8. The ordered pair solution is (171.28, 16.8)